# TEASMA: A Practical Approach for Test Assessment of Deep Neural Networks using Mutation Analysis

### v2

@@ Abstract
Successful deployment of Deep Neural Networks (DNNs), particularly in safety-critical systems, requires their validation with an adequate test set to ensure a sufficient degree of confidence in test outcomes. Mutation analysis, a well-established technique for measuring test adequacy in traditional software, has been adapted to DNNs in recent years. This technique is based on generating mutants that ideally aim to be representative of actual faults and thus can be used for test adequacy assessment. In this paper, we investigate for the first time whether and how mutation operators that directly modify the trained DNN model (i.e., post-training operators) can be used for reliably assessing the test inputs of DNNs. Our results show that these operators, though they do not aim to represent realistic faults, exhibit strong, non-linear relationships with faults. Inspired by this finding and considering the significant computational advantage of post-training operators compared to the operators that modify the training data or program (i.e., pre-training operators), we propose and evaluate TEASMA, an approach based on post-training mutation for assessing the adequacy of DNN’s test sets. In practice, TEASMA allows engineers to decide whether they will be able to trust test results and thus validate the DNN before its deployment. Based on a DNN model’s training set, TEASMA provides a methodology to build accurate DNN-specific prediction models of the Fault Detection Rate (FDR) of a test set from its mutation score, thus enabling its assessment. Our large empirical evaluation, across multiple DNN models, shows that predicted FDR values have a strong linear correlation ($R^{2}\geq 0.94$) with actual values. Consequently, empirical evidence suggests that TEASMA provides a reliable basis for confidently deciding whether to trust test results or improve the test set of a DNN model.
@@
Deep Neural Network, Test Assessment, Mutation Analysis.
@@ I Introduction
Effective testing of Deep Neural Networks (DNN) is essential to ensure they can be trusted, particularly as they are increasingly deployed in safety-critical systems. However, to rely on testing outcomes, it is crucial to assess the adequacy of the test set with which the DNN is to be validated. In the context of DNN testing, several techniques have been proposed for test adequacy assessment and are mostly relying on coverage criteria [1, 2, 3, 4, 5], inspired by practice in testing traditional software. However, these coverage criteria have shown no significant correlation with the number of mispredicted inputs or detected faults in a test set [6, 7, 5], thus putting into question whether they are reliable criteria for the test assessment of DNN models.
@@
Mutation analysis, one of the most used techniques for test assessment in traditional software [8, 9], has been adapted to DNN models in recent years. DNN-specific Mutation Operators (MO) have been proposed [10, 11, 12, 13] since traditional MOs are not directly applicable to DNNs [14]. Indeed, developing a DNN model entails applying a training process relying on both a training program and a training set. Therefore, researchers have proposed MOs that fall into two primary categories: post-training MOs [11, 10], which mutate the DNN model obtained after the training process, and pre-training MOs [11, 13], which mutate either the training program or the training set and use these mutated versions in the training process to generate mutated models.
Post-training MOs were introduced by mutation testing frameworks for DNN models: MuNN [10] and DeepMutation [11]. Although DeepMutation introduces both the pre- and post-training MOs, its authors later introduced DeepMutation++ [12], a mutation testing tool implementing only the post-training MOs. Post-training MOs offer a practical advantage by not requiring retraining for mutant generation, resulting in vastly reduced computational costs compared to pre-training MOs. However, they directly modify a DNN model and thus are criticized for not realistically representing real faults. Hembatova et al. [13] further expanded the concept of pre-training MOs by investigating real faults in DNN models [14] and proposing corresponding, more realistic pre-training MOs. They also introduced a statistical approach to determine whether a mutant generated by pre-training MOs is killed [14]. However, this approach requires generating many instances of the original DNN model and each mutated model by repeating the training process many times for each model. Such an expensive procedure makes the cost of mutation testing impractical in many situations.
The use of mutation analysis in traditional software is motivated by the strong association between mutants and faults reported in the literature [15, 16, 8]. Regarding DNNs, the ability of MOs to generate killable and non-trivial mutants has been thoroughly investigated [11, 13]. However, the association between the mutants and DNN faults, and thus the ability to assess a test set based on its Mutation Score (MS), has never been studied. In this paper, we address this gap by investigating such an association for mutants generated from post-training MOs, motivated by their much lower cost compared to pre-training MOs. For this purpose we investigate, across a number of models and input subsets from training sets, the relationship between MS and Fault Detection Rates (FDR) considering various MS definitions proposed in the literature [11, 12]. We perform our experiments using post-training MOs implemented by DeepMutation++, a State-of-The-Art (SoTA) mutation testing tool for Feed-Forward Neural Networks (FNNs). We show that there are strong, non-linear relationships between FDR and MS, suggesting that mutants generated by post-training MOs could be used as fault indicators.
We also show that, despite the strong relationships between FDR and MS, their shape is not only highly non-linear but widely varies across subjects. This observation indicates that, in the context of DNN models, MS calculated based on post-training MOs is not a surrogate measure for a test set’s fault detection capability and thus cannot be used directly for test assessment. These findings motivate us to propose a new methodology that leverages MS to build an accurate and specific FDR prediction model for each DNN. To this end, we propose TEASMA, a practical solution designed to assess the adequacy of test sets for DNN models before deployment through mutation analysis based on post-training MOs. Following TEASMA, engineers can develop a specific FDR prediction model, along with Prediction Intervals (PI), for each DNN model based on its training set. This prediction model can then be used by engineers to predict a test set’s FDR based on its MS and decide, considering the prediction’s PI, whether a test set is adequate for validation in their context. Developing TEASMA is further motivated by the costly and time-consuming process of preparing large test sets, whose labeling usually requires domain experts. Using TEASMA, a test set is assessed before labeling and execution, solely based on the training set used for training the DNN. Test inputs can then be labeled and executed once the test set is deemed adequate, thus preventing the waste of testing resources. Further, if sufficiently fast and when MS is shown to accurately predict FDR, TEASMA could be used to automatically guide test selection.
We evaluated TEASMA on five widely used DNN models and four image recognition input sets. We compared, for a large number of test subsets, their predicted FDR with the actual one and we could observe, in all cases, a very high linear correlation ($R^{2}\geq 0.94$ ) between the two, with a slope close to 1 for the regression line. We also show that our prediction models have narrow PIs, implying that the predicted FDR can confidently be used as a basis for deciding about the adequacy of a test set.
To further justify our choice of using post-training MOs in TEASMA, we replicated our experiments using pre-training MOs implemented by DeepCrime, considering only one of our smaller DNNs to maintain the computations at a reasonable level and obtain a conservative ballpark estimate of the cost of using pre-training MOs. Our results demonstrate that performing mutation analysis, using the same inputs based on pre-training MOs and the statistical MS calculation of DeepCrime, incurs a computational time approximately 153 times greater than that required for using DeepMutation++’s post-training MOs and MS calculation approach. To investigate whether such a substantially higher cost would be justified by a stronger association between mutants and faults, we compared the correlation between FDR and MS computed based on these two tools. The findings show that using MS based on pre-training MOs does not yield any advantage over post-training MOs. These results show that conducting mutation analysis using post-training MOs not only requires significantly less computational resources compared to pre-training MOs but also leads to a MS that can serve as a reliable basis for predicting FDR. This further confirms our choice of relying on post-training MOs in TEASMA.
@@
  • An empirical analysis of SoTA post-training MOs showing that these mutants have a strong relationship with faults in DNNs, though the shape of the relationship tends to be highly non-linear and varies widely across DNNs. Results thus confirm early suspicions that such MOs cannot directly be used as surrogates for faults and as such for test assessments.
  • TEASMA, a practical approach utilizing post-training MOs for assessing a test set before validation and deployment of a DNN. It is based on regression analysis using its training set, with the objective to build a predictor model for FDR, which can then be used as a basis for assessing the adequacy of test sets.
  • An empirical evaluation of TEASMA on five widely used DNN models and four image input sets, demonstrating its ability to accurately predict a test set’s FDR based on its MS, relying on a DNN-specific prediction model built during training.
  • An empirical analysis of the cost of applying pre-training MOs in comparison to post-training ones, to further justify our choice to rely on the latter.
The remainder of this paper is structured as follows: Section II provides background and discussions of the related work directly relevant to our research questions. Section III presents our proposed approach for test adequacy assessment. Section IV describes the experiments we performed to assess our approach. Section V presents and discusses the results for each research question. Section VI discusses the cost of pre-training MOs, and Section VII concludes the paper.
II Background and Related Work
This section is divided into four parts. In the first part, we briefly describe the proposed mutant generation processes for DNNs. The second part summarizes different ways of identifying killed mutants and calculating MS. In the third part, we discuss the association between mutants and faults and point out the challenge of investigating such association in DNNs. The last part describes a SoTA approach that can be leveraged to overcome this challenge and perform our experiments.
II-A Mutant Generation
@@
Humbatova et al. [13] proposed DeepCrime, a mutation testing tool that includes 24 pre-training mutation operators inspired by real faults in deep learning systems [17, 18, 19]. The main goal was to obtain more realistic mutation operators. They also proposed a statistical approach [14] for identifying killed mutants that requires multiple instances of each mutant. Therefore, DeepCrime repeats the training process on the original and modified versions of the training set and program many times, resulting in many instances of the original model and each mutant. This further increases the cost of mutation testing, requiring not only to repeat the training process and generating more mutants but also executing each test set on all mutant instances.
@@
II-B Mutation Score Calculation
Similar to traditional software, killed DNN mutants can be identified by comparing the output of a mutant $M_{i}$ with the original model $N$ and the corresponding MS can be calculated as:
@@ II-B Mutation Score Calculation
However, calculating MS as the average $KillingScore$ of test inputs in a test set may not indicate the ability of the entire test set to kill mutants. A test set cannot achieve such a high MS unless many of its test inputs are able to kill most of the mutants. This is visible in results reported by Humbatova et al. [13] where the MS calculated by $KSBasedMS$ for entire test sets ranges from 0.059 to 0.33. Humbatova et al. [13] further proposed a statistical approach for identifying killed mutants when applying pre-training MOs in DeepCrime where the training process is repeated $n$ times, leading to $n$ instances of the original model and mutants being generated (Section II-A). This further increases the cost of mutant generation using pre-training MOs and makes them computationally challenging for practical use in our context.
Humbatova et al. [13] proposed a statistical approach for identifying killed mutants when applying pre-training MOs in DeepCrime where the training process is repeated $n$ times and $n$ instances of the original model $N=\{N_{1},...,N_{n}\}$ and each mutant $M_{i}=\{M_{i1},...,M_{in}\}$ are generated (Section II-A). A mutant $M_{i}$ is considered killed by DeepCrime if, for a given test set $T$, the difference between the accuracy of $N(T)$ and $M_{i}(T)$ is statistically significant:
$isKilled(N,M_{i},T)=\begin{cases}1,&\text{if $p-value<\alpha\quad\textrm{and}\quad effectSize\geq\beta$}\\ 0,&\text{otherwise}\end{cases}$
The $p-value$ is obtained by using the Generalized Linear Model (GLM) [20] with $\alpha=0.05$ and $effectSize$ is calculated using Cohen’s d [21] with $\beta=0.5$. The MS of a test set can be calculated directly using $isKilled$ as follows:
$isKilledBasedMS(T,M)=\frac{\sum_{M_{i}\in M}isKilled(N,M_{i},T)}{\left|M\right|}$ (6)
DeepCrime [13], however, defined MS differently for a test set $T$ on a set of mutants generated by a given mutation operator $MO_{i}$:
$OperatorMS(MO_{i},T)=\frac{\left|K(MO_{i},T)\cap K(MO_{i},TrSet)\right|}{\left|K(MO_{i},TrSet)\right|}$
where $K(MO_{i},T)$ denotes those mutants generated by $MO_{i}$ that are killed by $T$, and $TrSet$, is the training input set used to generate the original model $N$. Only the intersection of mutants that are killed by both the test set and the training input set is considered in order to exclude equivalent mutants. Finally, the overall mutation score of $T$ on all mutants is determined as the average of mutation scores calculated for each individual operator:
$DeepCrimeMS(T,M)=\frac{\sum_{MO_{i}\in MO}OperatorMS(MO_{i},T)}{\left|MO\right|}$
This MS calculation captures the ability of a test set to kill mutants generated by all MOs and a test input subset can only achieve a high MS when it is able to kill a large proportion of mutants generated by each MO. It is also important to note that since the killed mutants are determined by comparing the accuracy of the original model and mutants based on Equation(5), the correct label of each test input must be known. Therefore, we cannot calculate the MS of an unlabeled test set before deployment with this approach.
To be able to use the pre-training MOs of DeepCrime and still calculate the MS of a test set before deployment, we can rely on comparing the output of a mutant with the original model and calculate MS in the same way as DeepMutation ($DeepMutationMS$ in Equation (2). This way we can also fairly compare the MS of a given test set on mutants generated by DeepCrime pre-training operators with DeepMutation post-training operators. For this purpose, we need to determine the output of each mutant and the original model. Since DeepCrime generates $n$ instances of the original model and each mutant, one way to determine the output is to rely on majority voting by considering the most voted output by all $n$ instances. Consequently, the output of a mutant $M_{i}$ (or the original model $N$) for a given test input $t$ can be identified as:
$M_{i}(t)=mode\{M_{i1}(t),M_{i2}(t),...,M_{in}(t)\}$ (9)
@@
II-C Mutation Analysis Cost
The high computational cost of mutation analysis is its main drawback in practice. This cost is mainly due to performing two processes. Initially, one must utilize MOs to generate mutants. The second process involves executing a large number of test inputs against mutants, followed by the calculation of their MS. A precise calculation of MS usually necessitates filtering out the equivalent, trivial, or redundant mutants.
The cost issue is even more critical in the case of DNNs considering large input sets, which is usual in practice. In this context, the substantial computational overhead of mutant generation using pre-training MOs, due to the need to perform training on each mutant, is a significant drawback when compared with post-training MOs. This is particularly true for DeepCrime which involves generating many instances of each mutant, even if only the most effective configurations [14] are considered for each MO to create non-equivalent and non-trivial mutants. Subsequently, computing MS is even more time-consuming as each input is run against all instances of each mutant. In order to characterize the differences in computational cost between mutation analysis employing pre- and post-training MOs, we report in Table I a formula capturing the average cost of each process as a function of parameters explained below. This is based on the assumption we use two SoTA mutation testing tools for DNNs, namely DeepMutation++ and DeepCrime.
TABLE I: Average cost function of performing mutation analysis using pre- or post-training MOs
Tool Mutant generation MS calculation
DeepMutation++ $mutateTime(N)\times\left|M_{G}\right|$ $execTime(T)\times(\left|M_{F}\right|+1)$
DeepCrime $mutateTime(TrSet,TrP)\times\left|M_{G}\right|$ $execTime(T)\times(\left|M_{F}\right|+1)\times n$
$+trainTime(M)\times\left|M_{G}\right|\times n$
First, $mutateTime(N)$ represents the average time to create a mutated version of the original model $N$. Similarly, $mutateTime(TrSet,TrP)$ denotes the average time to create a mutated version of either the original training set $TrSet$ or the training program $TrP$, the $trainTime(M)$ denotes the average time to train a DNN model, and the $execTime(T)$ represents the average time to execute a test set $T$ on either the original or a mutated DNN model. Additionally, $\left|M_{G}\right|$ is the number of generated mutants, and $\left|M_{F}\right|$ is the number of mutants after filtering out those considered equivalent, trivial, or redundant. The primary difference in mutant generation cost between any pre-training and post-training mutation testing tool stems from the retraining involved for each pre-training mutant. Given that DeepCrime generates $n$ instances for each mutant, the retraining process is performed $n$ times for every mutant. Consequently, the computational time required for calculating MS based on mutants generated by DeepCrime is $n$ times greater than that of DeepMutation++. Given the necessity of computing MS for numerous test sets during mutation analysis, the execution of these test sets results in a significant difference in computational cost between the two tools.
Considering such substantial differences, it is essential to investigate if the more practical post-training MOs are also able to generate mutants associated with faults that can support our objective, that is assessing the adequacy of test sets. In this paper, we perform such an investigation and propose TEASMA, a practical approach for test set adequacy assessment in DNNs based on mutation analysis using post-training MOs. Further, to support our analytical arguments above, we report on the cost of mutation analysis deploying pre-training MOs for one of our simplest DNNs and input sets in SectionVII. Results show that relying on DeepCrime’s MOs and MS calculation formula (Equation (8)) took about 153 times the amount of computation time of DeepMutation++. This ratio should vary across subjects and is logically expected to be higher for larger DNNs and input sets. Nevertheless, our observations confirm the expected high cost of using pre-training mutants.
Figure 2: The process of test adequacy assessment with TEASMA
II-D Mutants and Faults

### v3

# TEASMA: A Practical Methodology for Test Adequacy Assessment of Deep Neural Networks

@@ Abstract
Successful deployment of Deep Neural Networks (DNNs), particularly in safety-critical systems, requires their validation with an adequate test set to ensure a sufficient degree of confidence in test outcomes. Although well-established test adequacy assessment techniques from traditional software, such as mutation analysis and coverage criteria, have been adapted to DNNs in recent years, we still need to investigate their application within a comprehensive methodology for accurately predicting the fault detection ability of test sets and thus assessing their adequacy. In this paper, we propose and evaluate TEASMA, a comprehensive and practical methodology designed to accurately assess the adequacy of test sets for DNNs. In practice, TEASMA allows engineers to decide whether they can trust high-accuracy test results and thus validate the DNN before its deployment. Based on a DNN model’s training set, TEASMA provides a procedure to build accurate DNN-specific prediction models of the Fault Detection Rate (FDR) of a test set using an existing adequacy metric, thus enabling its assessment. We evaluated TEASMA with four state-of-the-art test adequacy metrics: Distance-based Surprise Coverage (DSC), Likelihood-based Surprise Coverage (LSC), Input Distribution Coverage (IDC), and Mutation Score (MS). We calculated MS based on mutation operators that directly modify the trained DNN model (i.e., post-training operators) due to their significant computational advantage compared to the operators that modify the DNN’s training set or program (i.e., pre-training operators). Our extensive empirical evaluation, conducted across multiple DNN models and input sets, including large input sets such as ImageNet, reveals a strong linear correlation between the predicted and actual FDR values derived from MS, DSC, and IDC, with minimum $R^{2}$ values of 0.94 for MS and 0.90 for DSC and IDC. Furthermore, a low average Root Mean Square Error (RMSE) of 9% between actual and predicted FDR values across all subjects, when relying on regression analysis and MS, demonstrates the latter’s superior accuracy when compared to DSC and IDC, with RMSE values of 0.17 and 0.18, respectively. Overall, these results suggest that TEASMA provides a reliable basis for confidently deciding whether to trust test results for DNN models.
@@
Deep Neural Network, Test Assessment, Test Adequacy Metrics.
@@ I Introduction
Effective testing of Deep Neural Networks (DNN) is essential to ensure they can be trusted, particularly as they are increasingly deployed in safety-critical systems. However, assuming the testing outcome is satisfactory, to be able to trust it, it is crucial to assess the adequacy of the test set with which the DNN is being validated. The goal of test adequacy assessment is therefore to ensure that a test set selected for evaluating a DNN model’s performance is effective at identifying faults and can thus confidently validate the model’s accuracy. In other words, when the DNN model achieves high accuracy on a test set, regardless of the test selection approach, the tester must determine if such results can be trusted and therefore whether to be confident about the model being deployed. Indeed, inadequate testing introduces the risk of unwarranted confidence in a model’s accuracy. More precisely, high-accuracy test results, on a test set with high adequacy (predicted FDR), indicate that the model has been rigorously evaluated and is likely to perform as expected on unseen inputs from the same input distribution. In this paper, we propose TEASMA (TEst ASsessment using Mutation Analysis), a practical and reliable methodology for accurately assessing the adequacy of unlabeled test sets. We thus expect that when a test set achieves high adequacy using TEASMA, the high-accuracy test outcomes based on that test set are deemed trustworthy indicators of a high-performance model.
@@
In the context of DNN testing, several techniques have been proposed for test adequacy assessment that are mainly relying on mutation analysis [1, 2, 3, 4] or coverage criteria [5, 6, 7, 8, 9, 10, 11]. Mutation analysis, one of the most used techniques for test assessment in traditional software [12, 13], has been adapted to DNN models in recent years. DNN-specific Mutation Operators (MO) have been proposed [1, 2, 3, 4] since traditional MOs are not directly applicable to DNNs [14]. Researchers have proposed MOs that fall into two primary categories: post-training MOs [2, 1], which mutate the DNN model obtained after the training process, and pre-training MOs [2, 4], which mutate either the training program or the training set and use these mutated artifacts in the training process to generate mutated models. Post-training MOs offer a practical advantage by not requiring retraining for mutant generation, resulting in vastly reduced computational costs compared to pre-training MOs. However, they directly modify a DNN model and thus are criticized for not realistically representing real faults [15]. The use of mutation analysis in traditional software is motivated by the strong association between mutants and faults reported in the literature [16, 17, 12]. However, the association between the mutants and DNN faults, and thus the ability to assess a test set based on its Mutation Score (MS), has never been studied. In this paper, we address this gap by investigating such an association for mutants generated by post-training MOs, motivated by their lower cost.
DNN testing has evolved to include a variety of coverage metrics inspired by coverage criteria for traditional software, including widely-used neuron and layer coverage criteria [8]. However, these criteria have shown no significant correlation with the number of mispredicted inputs or detected faults in a test set [18, 19, 20]. In addition to these coverage metrics, distribution-based coverage metrics including Surprise Coverage (SC) [7, 6] and Input Distribution Coverage (IDC) [5] have been introduced. The SC metrics proposed by Kim et al. [7, 6], including Distance-based SC (DSC) and Likelihood-based SC (LSC), distinguish themselves from previous coverage metrics by measuring how surprising test inputs are for the DNN model relative to the training inputs. Unlike neuron and layer coverage, which can often be satisfied with a few test inputs exhibiting common input patterns, SC measures the novelty of test inputs. IDC is a State-Of-The-Art (SOTA) coverage metric recently introduced by Dola et al. [5]. This innovative framework offers a novel black-box test adequacy metric specifically designed for DNNs. Unlike white-box coverage metrics such as SC, which assess the internal behavior of a DNN, IDC focuses on the input domain by employing a Variational Autoencoder (VAE) to learn a low-dimensional latent representation of the input distribution.
Though various test adequacy metrics have been proposed, such as MS, SC, and IDC, we need to devise and evaluate a comprehensive methodology for test adequacy assessment relying on such metrics. While these metrics have been validated to some extent, for example, the results of IDC demonstrating a correlation with misprediction, a practical test adequacy assessment methodology is still lacking. Further, these metrics need to be compared in this particular application context. Such a methodology can provide engineers with an accurate assessment of the ability of a test set to detect faults, enabling them to confidently decide whether the test set is adequate for validating the DNN.
In this paper, we address this gap by introducing and evaluating TEASMA, a comprehensive and practical methodology for test adequacy assessment. TEASMA relies on the training set and existing adequacy metrics to accurately predict the Fault Detection Rates (FDR) of a test set. Following TEASMA, engineers can develop a specific FDR prediction model using a selected adequacy metric, along with Prediction Intervals (PI), for each DNN model based on its training set. This prediction model can then be used by engineers to predict a test set’s FDR based on its Adequacy Score (AS), calculated using the selected metric. Considering the prediction’s PI, engineers can then decide whether a test set is adequate for validation in their context. While TEASMA provides the flexibility to leverage any adequacy metric, we apply TEASMA using four SOTA metrics, including DSC, LSC, IDC, and MS and perform a comprehensive investigation to determine which ones can predict a test set’s FDR accurately.
Developing TEASMA is further motivated by the costly and time-consuming process of preparing large test sets, as it allows for the assessment of test set adequacy without the need for human experts to label the test data, the latter being only required once the test set is considered adequate.
Using TEASMA, a test set is assessed before labeling and execution, solely based on the training set used for training the DNN. Test inputs can then be labeled and executed once the test set is deemed adequate, thus preventing the waste of testing resources.
We evaluated TEASMA using seven widely used DNN models and six image recognition input sets including ImageNet [21]. We first show that there are strong relationships between FDR and AS based on all investigated metrics, i.e., MS, DSC, LSC, and IDC. Specifically for MS, this suggests that mutants generated by post-training MOs could be used as fault indicators. However, despite these strong relationships, the shape of the relationships varies widely across subjects and is highly non-linear for most of them. Subsequently, we applied TEASMA, using all four metrics, to build DNN-specific FDR prediction models. We compared, for a large number of test subsets, their predicted FDR with the actual one. We could observe, for all metrics except LSC, a very high linear correlation between the two, with a slope close to 1 for the regression line. Consequently, we compared the accuracy of FDR predictions using MS, DSC, and IDC and concluded that MS is the most accurate metric for predicting FDR across all subjects.
@@
  • An empirical analysis of SOTA test adequacy metrics, including DSC, LSC, IDC, and MS, showing that AS calculated using these metrics has a strong relationship with FDR, though the shape of the relationship tends to be highly non-linear and varies widely across DNNs. Results thus confirm early suspicions that these metrics cannot directly be used as surrogates for FDR and thus for test assessments.
  • TEASMA, an accurate methodology utilizing existing SOTA test adequacy metrics for assessing a test set before the validation and deployment of a DNN. It is based on regression analysis using its training set, with the objective of building a predictor model for FDR, which can then be used as a basis for assessing the adequacy of test sets.
  • A large empirical evaluation of TEASMA, relying on four existing SOTA test adequacy metrics with seven widely used DNN models and six image input sets, that took approximately 40 days of runtime. Results suggest TEASMA has the ability to build accurate predictors of a test set’s FDR based on the DNN training set.
  • A comprehensive comparison between investigated metrics showing that MS calculated based on post-training MOs, is the most accurate metric for predicting FDR across all subjects, followed by DSC and IDC, whereas LSC proves to be ineffective for accurate FDR prediction.
The remainder of this paper is structured as follows: Section II provides background and discussions of the test adequacy assessment techniques for DNNs involved in our experiments. Section III presents our proposed test adequacy assessment methodology. Section IV describes the experiments we performed to assess our methodology. Section V presents and discusses the results for each research question. Section VI discusses the related work, and Section VII concludes the paper.
II Background
This section is divided into four parts. The first three parts provide a brief overview of existing SOTA test adequacy assessment techniques for DNNs involved in our experiments. These include mutation analysis and two distribution-based coverage metrics: the widely used Surprise Coverage (SC) and the recently introduced Input Distribution Coverage (IDC). We also discuss in detail the existing methods for calculating the Adequacy Score (AS) of a test set based on these metrics. The last part describes a SOTA approach that can be leveraged to investigate the association between AS of a test set based on these metrics and its fault detection ability, enabling us to perform our experiments.
II-A Mutation Analysis
In this section, we briefly overview mutation analysis for DNNs. We introduce two primary categories of existing Mutation Operators (MOs), namely pre- and post-training MOs, and discuss the cost of mutation analysis which is the key factor driving our choice for relying on post-training operators in TEASMA. Finally, we summarize different ways of calculating MS for post-training MOs.
II-A1 Mutant Generation
@@
Humbatova et al. [4] proposed DeepCrime, a mutation testing tool that includes 24 pre-training mutation operators inspired by real faults in deep learning systems [22, 23, 24]. The main goal was to obtain more realistic mutation operators. They also proposed a statistical approach [14] for identifying killed mutants that requires multiple instances of each mutant. Therefore, DeepCrime repeats the training process on the original and modified versions of the training set and program many times, resulting in many instances of the original model and each mutant. This further increases the cost of mutation testing, requiring not only to repeat the training process and generate more mutants but also executing each test set on all mutant instances.
@@
II-A2 Mutation Analysis Cost
The high computational cost of mutation analysis is its main drawback in practice. This cost is mainly due to performing two processes. Initially, one must utilize MOs to generate mutants. The second process involves executing a large number of test inputs against mutants, followed by the calculation of their MS. A precise calculation of MS usually necessitates filtering out the equivalent, trivial, or redundant mutants. The cost issue is even more critical in the case of DNNs considering large input sets, which is usual in practice. In this context, the substantial computational overhead of mutant generation using pre-training MOs, due to the need to perform training on each mutant, is a significant drawback when compared with post-training MOs.
In order to characterize the differences in computational cost between mutation analysis employing pre- and post-training MOs, we performed a preliminary study using pre-training mutants generated by DeepCrime with post-training mutants generated by DeepMutation++. Given the experiment’s time-intensive nature, which entails consuming enormous amounts of computing resources, we limited it to just one simple DNN and input set, specifically LeNet5 and MNIST (subject S1 introduced later in Table I). Conducting mutation analysis on the same set of input subsets randomly sampled from the MNIST training set took less than 49 minutes using DeepMutation++, compared to over 123 hours with DeepCrime–more than 150 times the duration required by DeepMutation++. Consequently, post-training MOs in DeepMutation++ require two orders of magnitude less computational resources than pre-training MOs. While this experiment, due to its high computational requirements, is conducted on a single subject of modest size, the findings underscore the substantial difference in computational resource requirements for mutation analysis based on pre- and post-training MOs and justify our choice to employ the latter.
Given the arguments and preliminary results presented above, we rely on the more practical post-training MOs in our experiments. We investigate if they are able to generate mutants associated with faults that can support our objective, which is assessing the adequacy of test sets. In this paper, we deploy post-training MOs in TEASMA for DNN test set adequacy assessment based on mutation analysis.
II-A3 Mutation Score Calculation
Various methods have been proposed to identify killed mutants and calculate MS based on post-training MOs. Similar to traditional software, killed DNN mutants can be identified by comparing the output of a mutant $M_{i}$ with the original model $N$ and the corresponding MS can be calculated as:
@@ II-A3 Mutation Score Calculation
However, calculating MS as the average $KillingScore$ of test inputs in a test set may not indicate the ability of the entire test set to kill mutants. A test set cannot achieve such a high MS unless many of its test inputs are able to kill most of the mutants. This is visible in results reported by Humbatova et al. [4] where the MS calculated by $KSBasedMS$ for entire test sets ranges from 0.059 to 0.33. Humbatova et al. [4] further proposed a statistical approach for identifying killed mutants when applying pre-training MOs in DeepCrime where the training process is repeated $n$ times, leading to $n$ instances of the original model and mutants being generated (Section II-A1). This further increases the cost of mutant generation using pre-training MOs and makes them computationally challenging for practical use in our context.
@@
II-A4 Mutants and Faults